The amount of money that Gali has to invest at a simple interest rate of 15% per year, to earn R5 250 interest after three years, is

[1] R5 249,48.

[2] R122 500,00.

[3] R11 666,67.

[4] R10 000,00

[5] R3 620,69.


1
Expert's answer
2021-06-15T17:50:44-0400

Given;

r=15%

SI=R5250

Time=3 years

"SI=P\times R\times T"

"R5250=P\times \frac{15}{100}\times 3"

"R5250=0.45P\\P=\frac{5250}{0.45}\\=R11,666.66667"


Gali should invest R11,666.66667


Hence the correct answer is [3] R11 666,67.